% notes for the August 21 meeting. Purpose: Status of Deployment. Attendees: Jielun Sun, Steve Oncley, Sean Burns, and Tony Delany. Some Items: ------------ * We think the ACD Balloon will be up and collecting data next week (they were planning to be up at the site today). * Jielun, Tony, Gordon, Andrew, Kurt, and Sean were out in the field yesterday (aug20). The thermocouples were connected to the CU data system and the Hydra was moved to the field location. Some new photos from yesterday are on the web page. * Hydra is in the field, but the LiCors are still in the Lab. Tony will get the LiCors finished today and tomorrow and then the final touches to Hydra will be finished on Monday of next week (Aug 26). The gas cylinders and the LiCors must be transported to the Hydra location. * Tram is to be setup outside of FL1 on Monday. The mini-towers have arrived and are to be setup at the site starting next week. The tram track should be marked with red/silver flaggine. * The IonOptics CO2 sensors have arrived, but they will probably need to be run with their own labview program and the data collected with an independent laptop system. * The ATD towers are collecting data now (though it's not avail from the webpage--i will check into this). * Brian Lamb is done and gone. He said they collected some good data and will supply some for the others.